5. Solve the following equation. Please show all of your work.
(1/7)^x = 343
can someone please help me with this problem?

Answer by ewatrrr(24785)

 
Hi, Re: domain of a composite fumction...Please Post on site , so others may benefit from the information.
Posting questions relating to e^x and e^(-x)
could also benefit many who use this site
(1/7)^x = 7^-x = 343 = (7)^3
-x = 3
x = -3
CHECKING our Answer***
(1/7)^(-3) = 1^-3/7^-3 = 7^3/1 = 343
Graph: y = 1-log(x) When x = 1, log(1) = 0 and y = 1 and when x = 10, y = 0
Plot Pt(1,1)& Pt(10,0) and use the y-axis as an asymptote
Might recommend the FREE graph software at http://www.padowan.dk.com